OpenStudy (anonymous):

My answer is B If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8 (4 points) Answer A. ) 4x + y = 8 -17y = 0 B.) 2x - 11y = 16 4x + 7y = -16 C.) -23y + 4z = 24 -17y = 0 D.) 13y - 2z = 24 7y - 4z = 0

OpenStudy (agent0smith):

So we want to isolate x in the second equation, which is x + 6y - z = -4 subtract 6y from both sides, add z to both sides to get: x = -4 - 6x + z Now substitute that into the first and third equation in place of x

OpenStudy (anonymous):

The answer I got was C...

OpenStudy (agent0smith):

oops, that should be x = -4 - 6y + z

OpenStudy (anonymous):

is C the wrong answer?

OpenStudy (agent0smith):

C looks correct :)

OpenStudy (anonymous):

thank you:))